# Can Input Attributions Interpret the Inductive Reasoning Process in In-Context Learning?

### v2

@@ 7 Related Work
Instance-based explanation seeks for the explanation in training data rather than the immediate input during inference as IA methods Wachter et al. (2017); Charpiat et al. (2019); Hanawa et al. (2021). These two paradigms of instance-based and IA-based explanations have been studied somewhat separately since the information source to seek the explanation is clearly different. On the other hand, in ICL, the training examples are now in the input during inference that can be analyzed by IA methods. In this sense, our investigation can seen as a new exploration of instance-based explanation with the help of IA methods.
@@ 8 Conclusions
We have pointed out and tackled the problem of interpreting the inductive reasoning process in ICL as a missing but reasonable milestone to be explored in LLM interpretability research. Our revisit to the IA methods in interpreting this ICL process has clarified their limitations from a new angle as well as provided fruitful insights and discussions on their practical usage in modern NLP. These findings have highlighted some issues in the community; in particular, even the fundamental task of mapping input and output has not been accomplished, and there is room to sophisticate previously developed interpretability tools to be suitable for LLMs.
@@
Our study has several limitations in scope. First, we focused primarily on popular gradient-based IA methods, leaving other approaches such as perturbation-based methods like LIME Ribeiro et al. (2016) and SHAP Lundberg and Lee (2017) for future work.
Regarding model selection, we concentrated on widely-used open-weight LLMs. Since applying IA methods requires gradient computation through backward propagation, computational constraints limited our ability to evaluate all available models, particularly large ones such as Llama-2-70B Touvron et al. (2023).
Our experimental design used synthetic tasks to better identify influential examples in the few-shot setting. While this approach allowed for controlled experimentation, both the number and format of tasks were limited. Future work could explore more realistic tasks with greater variations.
We focused exclusively on pre-training models, excluding post-training models from our analysis. This choice was motivated by our interest in basic few-shot learning, which is more commonly used with pre-trained models. Although post-training models might demonstrate higher accuracy on our tasks and the self-answer setting due to their potentially superior capabilities Riviere et al. (2024), our primary focus was on evaluating IA methods rather than model performance.
The optimization of the self-answer setting was not explored in depth, as our main interest lay in examining whether larger models showed improved performance in this setting rather than enhancing the setting itself.
Finally, while our ambiguous tasks were designed with two potential functions in mind, we acknowledge that models might interpret these tasks differently than intended. However, since our focus was on whether models could recognize these as ambiguous tasks with two possible answers and use specific examples to determine the appropriate response, we believe this limitation does not significantly impact our findings.
@@
This work advances our understanding of input attribution (IA) methods in the context of large language models’ (LLMs) in-context learning (ICL). Our findings contribute to the broader goal of developing more interpretable and safer AI systems by providing practical insights into the strengths and weaknesses of IA methods as tools for interpreting LLMs.
This study exclusively uses synthetic data generated through computational methods. No real user data, human annotations, or personally identifiable information was collected or used in our experiments. Our synthetic dataset generation process does not involve any human subjects, crowd workers, or demographic information.
Appendix A Finetuning details
@@ A.1 Finetuing parameters
A.2 Zero-shot task accuracy after finetuning
@@
Appendix B Base model results
@@
Appendix C Aggregating attritbuion with max
Appendix D Prompts
@@
Appendix E Chain-of-though format
@@
Appendix F Distribution of example with the highest attribution score
Figures 6, 7, 8, 9 and 10 present the distribution of example positions with the highest attribution scores. All IA methods, except for I $\times$ G, possess positional bias for certain tasks, specifically favoring examples either at the beginning or end, aligning with the position bias known to LLMs Liu et al. (2024).
Appendix G Informaiton for responsibility checklist
We utilized software libraries, including the Huggingface toolkit Wolf et al. (2020) (Apache License Version 2.0), Captum package Kokhlikyan et al. (2020) (BSD 3-Clause License) for IG computation, and PyTorch Ansel et al. (2024) (BSD 3-Clause License). These tools were used according to their licenses and intended usage. We used writing assistance tools (including Grammarly) for language error correction only. The computational budget of this work is approximately 600 GPU hours (with A100/H100/H200 machines).
@@ Appendix G Informaiton for responsibility checklist
Figure 6: Distribution of the positional of the example with the highest attribution scores across IA methods (Llama-2-7B model).
Figure 7: Distribution of the position of the example with the highest attribution scores across IA methods (Llama-2-13B model).
Figure 8: Distribution of the position of the example with the highest attribution scores across IA methods (Gemma-2-2B model).
Figure 9: Distribution of the position of the example with the highest attribution scores across IA methods (Gemma-2-9B model).
Figure 10: Distribution of the position of the example with the highest attribution scores across IA methods (Gemma-2-27B model).

### v5

@@ 7 Related Work
Instance-based explanation seeks explanation in the training data rather than relying on immediate input during inference, as IA methods do Wachter et al. (2017); Charpiat et al. (2019); Hanawa et al. (2021). These two paradigms of instance-based and IA-based explanations have been studied somewhat separately since the information source to seek the explanation is clearly different. On the other hand, in ICL, the training examples are now in the input during inference that can be analyzed by IA methods. In this sense, our investigation can be seen as a new exploration of instance-based explanation with the help of IA methods.
@@ 8 Conclusions
We have pointed out and tackled the problem of interpreting the inductive reasoning process in ICL as a missing but reasonable milestone to be explored in LLM interpretability research. Our revisit to the IA methods in interpreting the ICL process has clarified their limitations from a new angle as well as provided fruitful insights and discussions on their practical usage in modern NLP. Some methods frequently failed even on a very simple associative recall task, and there is room to sophisticate previously developed interpretability tools to be suitable for LLMs.
@@
Our study has several limitations in scope. First, we focused primarily on popular gradient-based IA methods, leaving other approaches such as perturbation-based methods like LIME Ribeiro et al. (2016) and SHAP Lundberg and Lee (2017) for future work. Furthermore, the optimization of the interpretation methods, e.g., the prompt used in the self-answer baseline, was not explored in depth.
Second, regarding model selection, we concentrated on widely-used open-weight LLMs. Since applying IA methods requires gradient computation through backward propagation, computational constraints limited our ability to evaluate all available models, particularly large ones such as Llama-2-70B Touvron et al. (2023). In addition, we focused exclusively on pre-trained models, excluding post-trained (e.g., instruction-tuned) models from our analysis, although this choice was motivated by starting from interpreting simpler few-shot abilities that emerged during the pre-training phase.
Third, our experimental design used synthetic tasks to better define influential examples in the few-shot setting. While this approach allowed for controlled experimentation, both the number and format of tasks were limited. The design of the synthetic dataset is also biased toward the input erasure method. Future work could explore more realistic tasks with greater variations.
Finally, we may have to acknowledge that models might interpret these tasks differently than intended. As described in the discussion, models may sometimes rely on their inductive biases that can not be attributed to input tokens. Tracing their rationale not only into input but also model internals may also be needed to fully interpret the model’s inductive reasoning process.
@@
This work explores input attribution (IA) methods over large language models’ (LLMs) in-context learning (ICL) ability. Our findings contribute to the broader goal of developing more interpretable and safer AI systems by providing practical insights into the strengths and weaknesses of IA methods as tools for interpreting LLMs. This study exclusively uses synthetic data generated through computational methods. No real user data, human annotations, or personally identifiable information were collected or used in our experiments. Our synthetic dataset generation process did not involve any human subjects, crowd workers, or demographic information. We employ LLMs for writing assistance, specifically for grammar checking and polishing the manuscript. All LLM outputs were subsequently reviewed and edited by the authors to ensure accuracy and fidelity to the intended message.
Acknowledgments
We want to express our gratitude to the members of the Tohoku NLP Group for their insightful comments. And special thanks to Sho Yokoi for his valuable suggestions on how to improve clarity in several aspects. This work was supported by the JSPS KAKENHI Grant Number JP24H00727, JP22J21492; JST Moonshot R&D Grant Number JPMJMS2011-35 (fundamental research); JST BOOST, Japan Grant Number JPMJBS2421; Grant-in-Aid for Early-Career Scientists Grant Number JP23K16938.
Appendix A Aggregating attritbuion with max
Appendix B Attribution with coalition
Figure 8 presents the IA scores obtained with coalition aggregation, that is, we treat each example as a single coalition by computing the mean pooling of all the token embeddings in the example to obtain its “coalition embedding”, and then calculate its attribution. Specifically, instead of using the embedding of each token $\mathbf{X}\coloneqq[x_{1},\ldots,x_{n}]$, we use the mean-pooled embedding $\bar{X}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$. The difference between calculating token-level attributions and then aggregating them with the coalition approach is minimal, except for the IxG metric in the Pos-Title task.
Appendix C Finetuning details
@@ C.1 Finetuing parameters
C.2 Zero-shot task accuracy after finetuning
@@
Appendix D Base model results
@@
Appendix E Prompts
@@
Appendix F Chain-of-though format
@@
Appendix G Distribution of example with the highest attribution score
Figures 9, 10, 11, 12, 13, 14 and 15 present the distribution of example positions with the highest attribution scores. All IA methods, except for I $\times$ G, possess positional bias for certain tasks, specifically favoring examples either at the beginning or end, aligning with the position bias known to LLMs Liu et al. (2024).
Appendix H Use of existing assets
Table 10 shows the assets being used in this paper, with the type, name, link, license, and citation for each asset used in the paper.
Appendix I Compute statement
Most experiments were conducted on a cluster equipped with NVIDIA A100 or H100 GPUs (80 GB each), while some fine-tuning runs were executed on H200 GPUs with 141 GB of memory. Fine-tuning time varies across models and tasks: the average run for the smallest model (Gemma-2-2B), takes roughly 2 hours per task, whereas the largest model (Gemma-2-27B), requires about 5 hours. IA evaluation time depends primarily on model size and the number of few-shot examples. Under the 10-shot setting, evaluation takes around 10 hours for (Gemma-2-2B) and about 24 hours for (Gemma-2-27B). In total, the experiments in this study consumed approximately 700 GPU hours, including exploratory runs.
@@ Appendix I Compute statement
Zero-shot (%) Few-shot (%)
Associative-Recall $0.5$ $100.0$ ($10.0$)
Linear-or-Distinct $50.0$ $99.8$ ($50.0$)
Add-or-Multiply $50.0$ $100.0$ ($50.0$)
Verb-Object $50.0$ $100.0$ ($50.0$)
Tense-Article $50.0$ $100.5$ ($50.0$)
Pos-Title $50.0$ $98.3$ ($50.0$)
Table 9: The zero-shot and few-shot accuracy of the fine-tuned Mistral-7B model across all evaluation tasks. The chance rate is indicated in parentheses.
@@
Figure 8: Coalition attribution accuracies for each task by treating each example as one coalition.
Figure 9: Distribution of the position of the example with the highest attribution scores across IA methods (Gemma-2-2B model), 50-Shot.
Figure 10: Distribution of the position of the example with the highest attribution scores across IA methods (Gemma-2-2B model, 100-Shot).
Figure 11: Distribution of the position of the example with the highest attribution scores across IA methods (Gemma-2-9B model, 10-Shot).
Figure 12: Distribution of the position of the example with the highest attribution scores across IA methods (Gemma-2-27B model, 10-Shot).
Figure 13: Distribution of the position of the example with the highest attribution scores across IA methods (Llama-2-7B model, 10-Shot).
Figure 14: Distribution of the position of the example with the highest attribution scores across IA methods (Llama-2-13B model, 10-Shot).
Figure 15: Distribution of the position of the example with the highest attribution scores across IA methods (Mistral-7B model, 10-Shot).
Asset Type Asset Name Link License Citation
[…]
Model Gemma-2-2B google/gemma-2-2b Gemma License Riviere et al. (2024)
Model Gemma-2-9B google/gemma-2-9b Gemma License Riviere et al. (2024)
Model Gemma-2-27B google/gemma-2-27b Gemma License Riviere et al. (2024)
Model Llama-2-7B meta-llama/Llama-2-7B Llama 2 Community License Touvron et al. (2023)
Model Llama-2-13B meta-llama/Llama-2-13B Llama 2 Community License Touvron et al. (2023)
Model Mistral-7B mistralai/Mistral-7B Apache license 2.0 Jiang et al. (2023)
Table 10: The list of assets used in this work.